# Adaptive Joint Estimation of Temporal Vertex and Edge Signals

### v2

@@ V-D COVID-19 case forecasting in population mobility network
For the England COVID-19 dataset, we will assume that the underlying topology is known when a new time instance arrives, but the simplicial signals on both the vertices and edges remain unknown and waiting to be predicted. During preprocessing, if multiple population mobility recordings exist between two vertices, we summed them up into a single edge and calculated the absolute sum of the population mobility. It should be emphasized that here population mobility is simply signals on the edges and should not be perceived using the concept of flows. The edge signals are scaled to $1/1000$ of the original value in the preprocessing and are scaled back to the original scale after obtaining the output. We added Gaussian noise to both the ground truth vertex and edge signals. The dynamic nature of this dataset is modeled by a dynamic sampling matrix in the edges $\mathbf{D}_{1,\mathcal{S}}[t]$ where we compare the graph topology of the graph at the current time point $\mathcal{G}[t]$ with newly obtained graph topology $\mathcal{G}[t+1]$. If an edge is present in both $\mathcal{G}[t]$ and $\mathcal{G}[t+1]$, then the corresponding diagonal element in $\mathbf{D}_{1,\mathcal{S}}[t+1]$ will be $1$. If an edge is present in $\mathcal{G}[t+1]$ but not in $\mathcal{G}[t]$, then the corresponding diagonal element in $\mathbf{D}_{1,\mathcal{S}}[t+1]$ will be 00. Both aggregations (17) and (18) are adopted in the experiment for predicting population mobility. The predicted number of confirmed cases by AJVEE will be compared against three online GSP algorithms: GLMS [4], GLMP [12], and G-Sign [11]. All these GSP algorithms are modified to be aware of the topology change, but they are not provided with the time-varying population mobility on the graph edges. The choice of this baseline is to demonstrate the usefulness of acquiring the population mobility at predicting the number of cases. We will also be using another simple baseline from [43] where the noisy cases of the previous day are used as the current estimation. This experiment reflects a real-world example of utilizing the time-varying signals on one simplicial dimension as an influence factor upon another simplicial dimension.
@@
The predicted population mobility and the predicted number of cases at three different time points using Algorithm 1 are illustrated in Fig. 7. From Fig. 7, it is clear that the recovered population mobility and the number of cases are close to the ground truth. In Fig. 8, we calculated the average absolute error of the predicted number of cases at each time point in all the vertices. By inspecting Fig. 8, we observe that our joint estimation framework has the lowest error most of the time. The results in this experiment indicate that our proposed AJVEE framework could utilize and predict jointly the dynamic information on the edges and the vertices to solve a real-world problem of online prediction of time-varying signals. The current approach to this real-world problem is a bit primitive. In the next section, we will discuss potential improvements to our AJVEE framework to solve similar problems.
@@
VI Conclusion and Future Work
@@
This paper presented the AJVEE framework for estimating time-varying vertex signals and edge signals, in which the dynamics of the vertex signals are in the form of time-varying edge signals. The underlying online estimation method of AJVEE is the ALMS-Hodge, an adaptive procedure backed by spatial and spectral TSP techniques. By treating graphs as a special case of a simplicial complex, two ALMS-Hodge with different simplicial orders are specified on the vertices and edges and then merged into the joint estimation. Even though AJVEE is proposed on the vertices and edges, we also discussed a more generalized setting where AJVEE is extended to higher-order structures. Experiments on both synthetic and real data indicate that AJVEE can efficiently solve the task of joint online estimation of time-varying vertex and edge signals.
@@
Currently, one limitation of AJVEE is the underlying topology is assumed to be known. One possible direction worth exploring is combining our AJVEE framework with topology learning algorithms. This combination will further extend the predictive power of AJVEE onto graphs that dynamically evolve the topology over time. In our current AJVEE framework, the influence from edge signals is being applied globally to all the signals on vertices using operations defined by the Hodge-Laplacian but without considering the change in directions. This leads to a limitation in our current setup that we are not able to process directed multi-edges between nodes. The extension of TSP approaches to multi-graphs and equivariant models will be another interesting direction.
Appendix: ALMS-Hodge Convergence analysis
@@
To make sure the ALMS-Hodge procedure in the AJVEE outputs a stable estimation, we analyze the update of the ALMS-Hodge algorithm under steady-state simplicial signal estimation. For ease of analysis, we will not be using the polynomial approximation of $\mathbf{H}_{k}$ shown in (9) in this analysis but rather use the spectral definition $\mathbf{H}_{k}=\mathbf{U}_{k,\mathcal{F}}(\mathbf{U}_{k,\mathcal{F}}^{T}\mathbf{D}_{k,\mathcal{S}}\mathbf{U}_{k,\mathcal{F}})^{-1}\mathbf{U}_{k,\mathcal{F}}^{T}$. Suppose that we have a clean simplicial signal $\boldsymbol{x}_{k}$, the update error $\tilde{\boldsymbol{x}}_{k}[t]$ at time $t$ is the difference between the current estimation $\hat{\boldsymbol{x}}_{k}[t]$ and $\boldsymbol{x}_{k}$: $\tilde{\boldsymbol{x}}_{k}[t]=\boldsymbol{x}_{k}-\hat{\boldsymbol{x}}_{k}[t]$. The ALMS-Hodge is stable if the update error converges under steady state estimation. To begin the proof, we use the SFT to transfer the update function (16) in the spectral domain:

### v3

@@ V-D COVID-19 case forecasting in population mobility network
For the England COVID-19 dataset, we will assume that the underlying topology is known when a new time instance arrives, but the simplicial signals on both the vertices and edges remain unknown and waiting to be predicted. During preprocessing, if multiple population mobility recordings exist between two vertices, we summed them up into a single edge and calculated the absolute sum of the population mobility. After the preprocessing, the population mobilities become signals on the edges and are not perceived using the concept of flows. The edge signals are scaled to $1/1000$ of the original value in the preprocessing and are scaled back to the original scale after obtaining the output. We added Gaussian noise to both the ground truth vertex and edge signals. The dynamic nature of this dataset is modeled by a dynamic sampling matrix in the edges $\mathbf{D}_{1,\mathcal{S}}[t]$ where we compare the graph topology of the graph at the current time point $\mathcal{G}[t]$ with newly obtained graph topology $\mathcal{G}[t+1]$. If an edge is present in both $\mathcal{G}[t]$ and $\mathcal{G}[t+1]$, then the corresponding diagonal element in $\mathbf{D}_{1,\mathcal{S}}[t+1]$ will be $1$. If an edge is present in $\mathcal{G}[t+1]$ but not in $\mathcal{G}[t]$, then the corresponding diagonal element in $\mathbf{D}_{1,\mathcal{S}}[t+1]$ will be 00. Both aggregations (16) and (17) are adopted in the experiment for predicting population mobility. The predicted number of confirmed cases by AJVEE will be compared against three online GSP algorithms: GLMS [4], GLMP [26], and G-Sign [29]. All these GSP algorithms are modified to be aware of the topology change, but they are not provided with the time-varying population mobility on the graph edges. The choice of this baseline is to demonstrate the usefulness of acquiring the population mobility at predicting the number of cases. We will also be using two simple baselines: 1. the MA of confirmed cases from the past 5 days; 2. repeat the noisy case count from the previous day (denoted as Last-day). This experiment reflects a real-world example of utilizing the time-varying signals on one simplicial dimension as an influence factor upon another simplicial dimension.
@@
The predicted population mobility and the predicted number of cases at three different time points using Algorithm 1 are illustrated in Fig. 7. From Fig. 7, it is clear that the recovered population mobility and the number of cases are close to the ground truth. In Fig. 8, we calculated the average absolute error of the predicted number of cases at each time point in all the vertices, which is the suggested metric by the dataset [36]. By inspecting Fig. 8, we observe that our joint estimation framework has the lowest error most of the time. AJVEE outperforms GSP algorithms that purely use vertex information because the GSP algorithms are unable to modify the magnitude of the update based on the information from the population mobility. However, because AJVEE jointly estimates the edge and vertex signals, AJVEE is more aware of the change in the vertices that are caused by the change in the edges. The nonstationary behavior of the confirmed cases causes relatively worse performance of MA and Last-day. The results in this experiment indicate that our proposed AJVEE framework could utilize and jointly predict the dynamic information on the edges and the vertices to solve a real-world problem of online prediction of time-varying signals. The current approach to this real-world problem is a bit primitive. In the next section, we will discuss potential improvements to our AJVEE framework to solve similar problems.
@@
VI Conclusion, Limitation, and Future Work
@@
This paper presented the AJVEE framework for estimating time-varying vertex signals and edge signals, in which the dynamics of the vertex signals are in the form of time-varying edge signals. The underlying online estimation method of AJVEE is the ALMS-Hodge, an adaptive procedure backed by spatial and spectral TSP techniques. By treating graphs as a special case of a simplicial complex, two ALMS-Hodge with different simplicial orders are specified on the vertices and edges and then merged into the joint estimation. Even though AJVEE is proposed on the vertices and edges, we also discussed a more generalized setting where AJVEE is extended to higher-order structures. In summary, AJVEE effectively addresses the challenges of representing topological irregularity, estimating unknown data from known data, removing noise present in the data, and extracting cross-space-time variations. Experiments on both synthetic and real data indicate that AJVEE can efficiently solve the task of joint online estimation of time-varying vertex and edge signals.
@@
Currently, one limitation of AJVEE is that the influence of edge signals is applied globally to all vertex signals using operations defined by the Hodge-Laplacian, without considering the direction. Even though the sign and directions at the time of edge estimation, when casting the edge signals as the influence of the vertex signals, the direction and sign information are not properly utilized, leading to an invariant model. However, these directional information and sign differences in edge signals, which can change over time, can be important to the data. The current invariant treatment of edge signals can lead to sub-optimal performance in downstream tasks that depend on these directional and sign attributes. We recognize this as a challenge and suggest that future work should explore equivariant models to better account for these variations, rather than relying on the current invariant model. Another limitation in our current setup is that we are not able to process directed multi-edges between nodes. Additionally, the underlying topology is assumed to be known. One possible direction worth exploring is combining our AJVEE framework with topology learning algorithms. This combination will further extend the predictive power of AJVEE onto graphs that dynamically evolve the topology over time. The extension of TSP approaches to dynamic graphs, multi-graphs, and equivariant models will be interesting future directions. Due to the lack of suitable real-world time-varying traffic flows for our experimental setting, we utilized a combination of synthetic data generated from real data on top of a graph representing real-world road networks. This is a limitation in the current experiment design and we look forward to applying our AJVEE framework to additional real-world datasets that contain real time-varying features.
Appendix A ALMS-Hodge Convergence analysis
@@
To make sure the ALMS-Hodge procedure in the AJVEE outputs a stable estimation, we analyze the update of the ALMS-Hodge under steady-state simplicial signal estimation. For ease of analysis, we will not be using the polynomial approximation of $\mathbf{H}_{k}$ shown in (27) in this analysis but rather use the spectral definition $\mathbf{H}_{k}=\mathbf{U}_{k,\mathcal{F}}(\mathbf{U}_{k,\mathcal{F}}^{T}\mathbf{D}_{k,\mathcal{S}}\mathbf{U}_{k,\mathcal{F}})^{-1}\mathbf{U}_{k,\mathcal{F}}^{T}$. Suppose that we have a clean simplicial signal $\boldsymbol{x}_{k}$, the update error $\tilde{\boldsymbol{x}}_{k}[t]$ at time $t$ is the difference between the current estimation $\hat{\boldsymbol{x}}_{k}[t]$ and $\boldsymbol{x}_{k}$: $\tilde{\boldsymbol{x}}_{k}[t]=\boldsymbol{x}_{k}-\hat{\boldsymbol{x}}_{k}[t]$. The ALMS-Hodge is stable if the update error converges under steady state estimation. To begin the proof, we use the SFT to transfer the update function (14) in the spectral domain:
@@
In order to observe the convergence behavior of the ALMS-Hodge, we will make online estimations for 1200 iterations where at each iteration we have a new noisy and missing observation of the steady-state edge signal. Gaussian distribution with mean $=0$ and variance $=0.18$ will be used for the noise distribution $\boldsymbol{\eta}[t]$. The step size is set to $\mu_{1}=0.1$. The low pass filter used in the experiment has the pass band of 26 lowest frequencies. Both spatial ALMS-Hodge with order P (denoted as Cheb P) and spectral ALMS-Hodge will be tested. We will be testing the ALMS-Hodge for 6 different settings: spectral ALMS-Hodge, Cheb 3, Cheb 5, Cheb 7, Cheb 9, and Cheb 11. The NMSE of this steady-state experiment is shown in Fig. 9.
Figure 9: Steady-state estimation on the edges of the Sioux Falls network using ALMS-Hodge with different step sizes.
By inspecting Fig 9, we can observe that all variants of the ALMS-Hodge result in similar NMSE results after 200 iterations. Additionally, all the spatial update functions have similar NMSE performance to the spectral update function when converged.
Appendix B Spectral analysis of AJVEE